Hastings Corporation is interested in acquiring Vandell Corporation. Vandell has 1 million shares outstanding and a target capital structure consisting of 30% debt; its beta is 1.30 (given its target capital structure). Vandell has $8.35 million in debt that trades at par and pays an 7.6% interest rate. Vandell’s free cash flow (FCF0) is $2 million per year and is expected to grow at a constant rate of 6% a year. Vandell pays a 35% combined federal and state tax rate. The risk-free rate of interest is 3% and the market risk premium is 4%. Hastings’ first step is to estimate the current intrinsic value of Vandell.

A) What are Vandell’s cost of equity and weighted average cost of capital? Round your answer to two decimal places. Do not round intermediate calculations.

Cost of equity: ___%

WACC: _____%

B) What is Vandell's intrinsic value of operations? (Hint: Use the free cash flow corporate valuation model.) Round your answer to two decimal places. Do not round intermediate calculations.

$_______ million

C) What is the current intrinsic value of Vandell's stock? Round your answer to the nearest cent. Do not round intermediate calculations.

$______ /share
Answer #1
A) Cost of equity as per CAPM = 3%+1.30*4% = 8.20%
After tax cost of debt = 7.6%*(1-35%) = 4.94%
WACC = 8.20%*70%+4.94%*30% = 7.22%
B) Intrinsic value of operations = 2*1.06/(0.0722-0.06) = $     173.77 million
C) Intrinsic value of equity = 173.77 - 35.00 = $     138.77 million
# of shares outstanding 1.00 million
Current intrinsic value of stock $     138.77